# ode]python

### v1

@@ 2.3 Capacitance model
$E_{ij}=\frac{1}{Z_{i}Z_{j}}\left[c_{k}\delta_{ij}\int_{a_{i}}^{b_{i}}n(x)^{2}dx+\frac{1}{2}\int_{a_{i}}^{b_{i}}\int_{a_{j}}^{b_{j}}K(x,x^{\prime})n(x)n(x^{\prime})dxdx^{\prime}\right]$ (15)
@@ 3 Data generation
QDFlow contains convenience functions for generating CSDs and rays in the generate module. Since the primary purpose of QDFlow is to generate data for training specialized ML models, it is essential that the resulting dataset captures the full range of variability observed in contemporary QD devices. To achieve this, QDFlow includes functionality to randomize nearly all physics parameters and to control the distributions from which each parameter is drawn. This capability is implemented via the PhysicsRandomization dataclass, which specifies each physics parameter as either a fixed value (when no randomization is desired, e.g., to allow regeneration of the same QD device), or a Distribution from which to draw the random values. A code example in Listing 3 shows how to import QDFlow and how to initialize a random configuration of physical parameters, with $\mu$ drawn from a distribution provided by the user.
@@
While CSDs are extremely useful for visualizing charge states, they require extensive data collection. In practice, only a relatively small subset of points–the charge transitions–are of primary importance. As the number of dots grows, so does the dimensionality of the gate voltage space, rapidly making the exploration of the complete, multidimensional voltage space infeasible. This challenge is typically handled by measuring multiple 2D CSDs, each defined by a different pair of gates.
@@
from qdflow import generate import numpy as np import itertools
@@
# Generate quasirandom points inside a given area num_points = 50 point_generator = generate.quasirandom_points(dim=2, min=2., max=16.) initial_points = itertools.islice(point_generator, num_points)
@@ 4 Noise
Figure 7(a) shows an example of a noiseless CSD simulated with QDFlow. CSDs with noise implementations adapted from QFlow—white, pink, and telegraph noise and Coulomb peak—are depicted in panels (b), (c), (d), and (e), respectively. CSDs with latching, $\operatorname{sech}^{2}$ blur, unintended QD, and sensor-gate coupling—new to the QDFlow—are presented in panels (f), (g), (h), and (i), respectively.
The simplest, white noise, is implemented by adding to each pixel a value drawn from a normal distribution with standard deviation given by the magnitude of the white noise. Pink noise is implemented by generating white noise in Fourier space with a uniform random phase and magnitude proportional to $1/\sqrt{f_{x}^{2}+f_{y}^{2}}$, where $f_{x}$ and $f_{y}$ are the components of each point in Fourier space, and then applying an inverse Fourier transform.
@@
The latching noise implemented in QDFlow can be controlled using one of two methods. The first and simpler legacy method to simulate latching is to shift each row of pixels by a random number of pixels drawn from a geometric distribution. This process produces a latching-like effect along the charge transitions; however, it is somewhat unrealistic since all transitions on the same row are shifted by the same amount, and because pixels far away from transitions are also displaced. A more physically realistic method relies on the nanowire simulation to calculate the sensor readout for both an excited charge state and a stable state at each pixel. The excited state chosen corresponds to the charge configuration most recently occupied prior to the most recent transition when sweeping the gate voltages. When latching noise is added, the sensor readout from the excited state replaces the stable-state readout for the first few pixels after each transition, with the number of pixels randomized each instance. In general, the second method is preferable as it more accurately reflects the experimental conditions; however, the first, legacy method is provided as a fallback when excited-state data are unavailable or computationally inconvenient, or impossible to obtain.
@@
Given the computational complexity of Thomas–Fermi calculations, the noise module is configured to assume that a complete noiseless CSD (or a ray-based data) has already been generated using the generate module. This approach gives us several advantages. First, it significantly reduces the computational overhead since multiple noise realizations with different relative noise strengths can be generated from a single noiseless CSD. Secondly, the modular approach adopted in QDFlow provides flexibility that cannot be achieved experimentally, where it is not possible to calibrate individual noise sources to the desired level. Additionally, it allows us to specify the magnitudes of each of the noise types relative to the local scale of the surrounding data points, which is important since the scale of the data points can vary across large CSDs. Finally, it allows for systematic benchmarking of ML algorithms since the noise level and type can be modified independently of the underlying physical configuration. For example, one can generate a single noiseless dataset and then apply different realizations of noise to study the robustness of a given algorithm under varying experimental conditions, as was done in Ref. 12.
@@ 5 Conclusion
Progress toward scalable quantum information technologies based on QD systems depends critically on overcoming the complexity of device operation and calibration as arrays increase in size. Novel ML-based methods have emerged as powerful tools to address these challenges, but their effectiveness relies on access to large, diverse, and realistically labeled datasets. QDFlow was developed precisely to meet this need.
QDFlow differs from existing quantum dot simulations in that it fully simulates the charge density function $n(x)$. By integrating a self-consistent Thomas–Fermi solver with a dynamic capacitance model, QDFlow provides a physics-informed simulation framework that goes beyond constant-capacitance approximations. This enables the generation of CSDs and ray-based data with features that evolve naturally with gate voltages, mimicking experimental behavior such as dot merging and transition slope variation.
The modular data generation tools allow for extensive randomization over physical parameters, yielding highly diverse synthetic datasets suitable for ML applications, while the noise module introduces experimentally relevant effects—including thermal broadening, telegraph noise, latching, and unintended transitions—in a controllable fashion. Together, these features make QDFlow uniquely positioned to support both the development and benchmarking of ML algorithms implemented in a wide range of tuning procedures. Early use cases, such as the QFlow-lite and QFlow 2.0 datasets, have already demonstrated QDFlow’s ability to accelerate the training of ML models for global state recognition [19, 8], ray-based navigation and charge tuning [11, 12], data quality assessment [12], detection of spurious QDs [29], and virtualization of QD arrays[13]. As an open-source, extensible platform with comprehensive documentation, QDFlow is designed to serve as both a research tool and a community resource.

### v2

@@ 2.3 Capacitance model
$E_{ij}=\frac{1}{Z_{i}Z_{j}}\left[c_{k}{}_{ij}\int_{a_{i}}^{b_{i}}n(x)^{2}dx+\frac{1}{2}\int_{a_{i}}^{b_{i}}\int_{a_{j}}^{b_{j}}K(x,x^{\prime})n(x)n(x^{\prime})dxdx^{\prime}\right]$ (15)
@@ 3 Data generation
QDFlow contains convenience functions for generating CSDs and rays in the generate module. Since the primary purpose of QDFlow is to generate data for training specialized ML models, it is essential that the resulting dataset captures the full range of variability observed in contemporary QD devices. To achieve this, QDFlow includes functionality to randomize nearly all physics parameters and to control the distributions from which each parameter is drawn. This capability is implemented via the PhysicsRandomization dataclass, which specifies each physics parameter as either a fixed value (when no randomization is desired, e.g., to allow regeneration of the same QD device), or a Distribution from which to draw the random values. A code example in Listing 3 shows how to import QDFlow and how to initialize a random configuration of physical parameters, with drawn from a distribution provided by the user.
@@
While CSDs are extremely useful for visualizing charge states, they require extensive data collection. In practice, only a relatively small subset of points—the charge transitions—are of primary importance. As the number of dots grows, so does the dimensionality of the gate voltage space, rapidly making the exploration of the complete, multidimensional voltage space infeasible. This challenge is typically handled by measuring multiple 2D CSDs, each defined by a different pair of gates.
@@
from qdflow import generate import numpy as np from scipy.stats import qmc
@@
# Generate quasirandom points inside a given area v_min, v_max = 2., 16. point_generator = qmc.Halton(d=2, scramble=False) initial_points = qmc.scale(point_generator.random(n=50), v_min, v_max)
@@ 4 Noise
Figure 7(a) shows an example of a noiseless CSD simulated with QDFlow. CSDs with noise implementations adapted from QFlow—white, pink, and telegraph noise and Coulomb peak—are depicted in panels (b), (c), (d), and (e), respectively. CSDs with latching, $\operatorname{sech}^{2}$ blur, unintended QD, and sensor-gate coupling—new to QDFlow—are presented in panels (f), (g), (h), and (i), respectively.
The simplest, white noise, is implemented by adding to each pixel a value drawn from a normal distribution with standard deviation given by the magnitude of the white noise. Pink noise is implemented by generating white noise in Fourier space with a uniform random phase and magnitude proportional to $1/\sqrt{f_{x}^{2}+f_{y}^{2}}$, where $f_{x}$ and $f_{y}$ are the components of each point in Fourier space, and then applying an inverse Fourier transform. This configuration allows for greater variability compared to telegraph noise. Alternatively, QDFlow also provides the option to add pink noise correlated along only the primary measurement axis, which corresponds to the more physical model of pink noise correlated in time.
@@
The latching noise implemented in QDFlow can be controlled using one of two methods. The first and simpler legacy method to simulate latching is to shift each row of pixels by a random number of pixels drawn from a geometric distribution. This process produces a latching-like effect along the charge transitions; however, it is somewhat unrealistic since all transitions on the same row are shifted by the same amount, and because pixels far away from transitions are also displaced. A more physically realistic method relies on the nanowire simulation to calculate the sensor readout for both an excited charge state and a stable state at each pixel, similar to what is implemented in Ref. 17. The excited state chosen corresponds to the charge configuration most recently occupied prior to the most recent transition when sweeping the gate voltages. When latching noise is added, the sensor readout from the excited state replaces the stable-state readout for the first few pixels after each transition, with the number of pixels randomized each instance. In general, the second method is preferable as it more accurately reflects the experimental conditions; however, the first, legacy method is provided as a fallback when excited-state data are unavailable or computationally inconvenient, or impossible to obtain.
@@
Given the computational complexity of Thomas-Fermi calculations, the noise module is configured to assume that a complete noiseless CSD (or a ray-based data) has already been generated using the generate module. This approach gives us several advantages. First, it significantly reduces the computational overhead since multiple noise realizations with different relative noise strengths can be generated from a single noiseless CSD. Secondly, the modular approach adopted in QDFlow provides flexibility that cannot be achieved experimentally, where it is not possible to calibrate individual noise sources to the desired level. Additionally, it allows us to specify the magnitudes of each of the noise types relative to the local scale of the surrounding data points, which is important since the scale of the data points can vary across large CSDs. Finally, it allows for systematic benchmarking of ML algorithms since the noise level and type can be modified independently of the underlying physical configuration. For example, one can generate a single noiseless dataset and then apply different realizations of noise to study the robustness of a given algorithm under varying experimental conditions, as was done in Ref. 12.
@@ 5 Conclusion
Progress toward scalable quantum information technologies based on QD systems depends critically on overcoming the complexity of device operation and calibration with increasing number of QDs. Novel ML-based methods have emerged as powerful tools to address these challenges, but their effectiveness relies on access to large, diverse, and accurately labeled datasets. QDFlow was developed precisely to meet this need.
QDFlow differs from existing QD simulations in that it fully simulates the charge density function $n(x)$. By integrating a self-consistent Thomas-Fermi solver with a dynamic capacitance model, QDFlow provides a physics-informed simulation framework that goes beyond constant-capacitance approximations. This enables the generation of CSDs and ray-based data with features that evolve naturally with gate voltages, mimicking experimental behavior such as dot merging and transition slope variation.
The modular data generation tools allow for extensive randomization over physical parameters, yielding highly diverse synthetic datasets suitable for ML applications, while the noise module introduces experimentally relevant effects—including thermal broadening, telegraph noise, latching, and unintended transitions—in a controllable fashion. Together, these features make QDFlow uniquely positioned to support both the development and benchmarking of ML algorithms implemented in a wide range of tuning procedures, device architectures, and material platforms. Early use cases, such as the QFlow-lite and QFlow 2.0 datasets, have already demonstrated QDFlow’s ability to accelerate the training of ML models for global state recognition [19, 8], ray-based navigation and charge tuning [11, 12], data quality assessment [12], detection of spurious QDs [29], and virtualization of QD arrays[12, 13]. As an open-source, extensible platform with comprehensive documentation, QDFlow is designed to serve as both a research tool and a community resource.
@@ Acknowledgements
Code availability
QDFlow is available on the Python Package Index, with the source code released under the GNU General Public License, and can be installed using pip install QDFlow command. The associated GitHub repository is https://github.com/QDFlow/QDFlow-solver. Any discovered bugs should be reported using GitHub issues. If you find this package useful, please star the repository and cite this paper.